# Zhiyuan Fan     Yumeng Wang     Sandeep Polisetty     Yi R. (May) Fung

### v2

@@ Abstract
Large Vision Language Models (LVLMs) excel in various vision-language tasks. Yet, their robustness to visual variations in position, scale, orientation, and context that objects in natural scenes inevitably exhibit due to changes in viewpoint and environment remains largely underexplored. To bridge this gap, we introduce V2R-Bench, a comprehensive benchmark framework for evaluating Visual Variation Robustness of LVLMs, which encompasses automated evaluation dataset generation and principled metrics for thorough robustness assessment. Through extensive evaluation on 21 LVLMs, we reveal a surprising vulnerability to visual variations, in which even advanced models that excel at complex vision-language tasks significantly underperform on simple tasks such as object recognition. Interestingly, these models exhibit a distinct visual position bias that contradicts theories of effective receptive fields, and demonstrate a human-like visual acuity threshold. To identify the source of these vulnerabilities, we present a systematic framework for component-level analysis, featuring a novel visualization approach for aligned visual features. Results show that these vulnerabilities stem from error accumulation in the pipeline architecture and inadequate multimodal alignment. Complementary experiments with synthetic data further demonstrate that these limitations are fundamentally architectural deficiencies, scoring the need for architectural innovations in future LVLM designs. [^1] Our code and data will be available at https://github.com/toward-agi/Visual-Variations-Robustness.
[Uncaptioned image]  V2R-Bench: Holistically Evaluating LVLM Robustness to Fundamental Visual Variations
Zhiyuan Fanthanks: Equal contribution.     Yumeng Wang     Sandeep Polisetty     Yi R. (May) Fung Hong Kong University of Science and Technology zhiyuan.fan@connect.ust.hk
Figure 1: Illustration of V2R-Bench. Left: Automated data generation framework that adds visual variations to images, covering basic visual tasks, existing benchmarks, as well as supplementary visual tasks. Right: Metrics for robustness testing, including performance consistency, semantic and token-level stability, and LLM-as-a-judge.
@@ 1 Introduction
While current benchmarks extensively evaluate models using images collected in specific scenarios, they overlook a more fundamental generalization capability: the robustness of LVLMs to visual variations. In input images, objects naturally exhibit diverse variations: spatial positions shift with changes in camera angles and viewpoints; object scales vary depending on viewing distances; orientations deviate from standard poses through rotations and inversions; and objects appear in a range of visual semantic contexts. These visual variations raise concerns about LVLMs’ robustness: whether models maintain consistent perception capabilities across all spatial positions in input images; what visual acuity threshold determines reliable performance; and how changes in orientation and visual context influence model behavior.
Despite the importance of such robustness, current research has primarily focused on model robustness to socio-cultural factors (Ananthram et al., 2024), adversarial prompt attacks (Wu et al., 2024b; Liu et al., 2024b), or corrupted image inputs (Liu et al., 2024a), leaving the impact of natural visual variations largely unexplored.
@@ 2 Related Work
Large Vision Language Models.
The remarkable progress of LVLMs stems from continuous advances in model architectures and the exponential growth of training datasets. On the architectural front, the rapid progress in foundation language models (OpenAI et al., 2024; Touvron et al., 2023; Jiang et al., 2023; Bai et al., 2023a) and vision models (Radford et al., 2021; Oquab et al., 2024), together with increasingly sophisticated multimodal alignment modules (Liu et al., 2023; Li et al., 2023b; Alayrac et al., 2022; Zhu et al., 2023; Wang et al., 2024b; Zhou et al., 2024), has established the fundamental capabilities of LVLMs in understanding and reasoning across modalities. Building upon these architectural foundations, the emergence of high-quality datasets (Chen et al., 2023a; Zhao et al., 2023; Wang et al., 2023b; Li et al., 2023c) designed for different training stages (e.g., multimodal alignment, visual instruction tuning, preference alignment) has enabled LVLMs to demonstrate exceptional performance across diverse real-world scenarios. More recently, research efforts have focused on optimizing lightweight model architectures and curating datasets tailored to edge scenarios (Ma et al., 2024; Chu et al., 2024; Zhu et al., 2024; Yao et al., 2024; Lee et al., 2025), thereby promoting the practical deployment and adoption of LVLMs in real-world applications.
While the pipeline architecture of these LVLMs effectively leverages pretrained domain-specific knowledge from vision and language components, it potentially accumulates errors and vulnerabilities across modules. However, no systematic investigation has been conducted to attribute model failures to individual architectural components, limiting both model interpretability and understanding of modality alignment.
@@
3 Evaluation Framework
In this section, we present our V2R-Bench evaluation framework, which includes an automated pipeline for generating diverse visual variations across task settings (3.1), our evaluation protocols (3.2), and a detailed description of the construction and characteristics of our evaluation dataset (3.3).
@@ 3.1 Automated Variations Generation
The automated data generation pipeline incorporates four fundamental visual variations that are prevalent in real-world scenarios. Position variations investigate whether LVLMs exhibit blind spots in their visual processing, where models may fail to accurately perceive visual information at certain spatial locations. Meanwhile, scale variations examine the perceptual boundaries of LVLMs when processing objects at different scales, similar to clinical vision tests for humans. Orientation variations challenge the ability of LVLMs to process objects at different rotational angles, which is crucial for real-world scenarios like robotic navigation where objects rarely appear in canonical orientations. Moreover, context variations test whether model predictions remain consistent across diverse environmental settings, to reveal whether LVLMs perform genuine visual perception or rely primarily on contextual cues for inference.
@@
where $P=\{1,...,W\}\times\{1,...,H\}$, $S=[s_{min},s_{max}]$, $R=[0,2\pi]$, and $C\in\mathcal{B}$ represent the sets of position, scale, rotation, and context variations respectively. For each original question-image pair, this generation process produces $|P|\times|S|\times|R|\times|C|$ variants in total, enabling a holistic exploration of the entire variation space that leaves no potential vulnerability unexplored.
@@ 3.2 Evaluation Metrics
Furthermore, LLM-as-a-judge (Zheng et al., 2023) is employed to emulate human assessment of LVLM-generated outputs under structured visual variations, providing an additional qualitative perspective on model robustness.
3.3 Dataset Construction
The proposed automated generation pipeline is implemented across two categories of tasks. The first category examines fundamental visual capabilities through object recognition and direction recognition tasks, with target objects and directional indicators being systematically transformed through our visual variations using image processing algorithms and inpainting diffusion models (Corneanu et al., 2024; Lugmayr et al., 2022)
The second category extends to existing multimodal benchmarks, focusing on scenarios where variations in position, scale, orientation, and context preserve ground-truth validity, ensuring that any performance changes reflect model robustness rather than ground-truth alteration. The final evaluation datasets contain a total of 428K images. Each category serves a distinct yet complementary purpose in our evaluation: the basic tasks provide controlled, interpretable measures of fundamental capabilities, while the extended benchmarks assess robustness in more naturalistic settings. The detailed implementation of these generation algorithms is provided in Appendix C.
@@ 3.3 Dataset Construction
Table 1: Comprehensive evaluation of model performance (i.e., accuracy and robustness) under different visual variations, for object and direction detection tasks with metrics defined in Section 3.2.
@@ 4.1 Component-level Analysis
$v=E_{v}(I;\theta_{v})\in\mathbb{R}^{N_{v}\times D_{v}}\vspace{-0.4em}$ (5)
@@
$h=P(v;\theta_{p})\in\mathbb{R}^{N_{v}\times D_{h}}\vspace{-0.4em}$ (6)
@@
The multimodal projector $P$ acts as a bridge between the visual feature and language embedding spaces, which raises two questions for its performance analysis: (RQ1) Do the projected features $h$ preserve the visual information contained in $v$?, and (RQ2) Do the projected features $h$ align well with the embedding space of the language model $M$? To answer these questions, two analytical approaches are employed: (1) comparing the performance of pre-projection features $v$ and post-projection features $h$ on visual tasks to quantify potential visual information loss during multimodal alignment, and (2) measuring the discriminability between projected visual features $h$ and the language embeddings of their corresponding captions to assess the quality of modality alignment.
@@ 4.2 Visual-Linguistic Feature Analysis
$t=\text{topk}\left(\text{softmax}\left(hE^{\top}\right)\right),\vspace{-0.55em}$ (8)
@@
Through these decoded language tokens, the semantic meaning captured by the aligned visual features becomes intuitively understandable, which helps address the following questions: (RQ3) how do language models interpret these aligned visual features? (RQ4) How robust is the semantics of aligned visual features to visual variations? These questions are investigated in detail in Section  5.2.
@@ 5.1 Impact of Visual Variations
The evaluation results across different LVLM archi-
tectures and model scales are presented in Tables 1. Despite their impressive performance demonstrated on complex visual tasks, these models exhibit surprising vulnerability to simple visual variations, resulting in significantly degraded performance across basic visual tasks. Even proprietary models such as gpt-4o, Claude, and Gemini not only produce incorrect outputs, but also occasionally claim an inability to perceive the visual content entirely, as illustrated in Appendix F. Interestingly, despite claims that some carefully distilled smaller models outperform their larger counterparts on mainstream benchmarks, our analysis reveals that scaling laws still hold for robustness: within the same model architecture, larger models consistently demonstrate better stability across visual variations.
@@
Table 2: Quantitative analysis of position bias. Experimental results show that these models consistently demonstrate better perception of surrounding regions compared to the central areas.
@@ 5.2 Component Analysis
Table 4: Linear probing accuracy (%) of vision language models on Object (Obj.) and Direction (Dir.) tasks reveals the collaborative impact of vision encoders/multimodal projectors and LVLMs. Notably, Qwen2-VL, which employs a native vision encoder trained directly with the language model rather than pretraining on image modality first and then aligning to language modality, shows no performance loss and even improvement after modality projection, suggesting the potential of unified architectures.
@@
Table 5: Performance of LVLM language model backbones on text-based tasks: comparing scenarios with tokenized background represented as asterisks (w/ BG) versus random words (w/o BG).
@@ 7 Conclusion
We present V2R-Bench, an evaluation framework designed to assess the robustness of LVLMs against visual variations. Our results show significant vulnerabilities in existing LVLMs and identify their origins as twofold: insufficient multimodal alignment and error accumulation inherent in pipeline model architectures. While synthetic data augmentation Qin et al. (2025) seems as a mitigation strategy, fundamental advancement in the field requires a shift towards native multimodal architectures rather than the current approach of concatenating separate language and vision modalities. We aim to draw greater attention to the importance of visual robustness in LVLMs and inspire future research toward more robust architectural paradigms.
@@
Aligned with the paper track’s focus, this work primarily concentrates on identifying a novel problem, establishing corresponding evaluation methods, and providing initial analytical insights into these vulnerabilities. Regarding solutions to the identified vulnerabilities, only two data synthesis approaches were explored, while potential improvements through architectural modifications, pre-training strategies, test-time self-correction or self-consistency approaches He et al. (2024); Wang et al. (2025), and novel modality alignment methods remain unexplored due to their substantial resource requirements. The detailed analysis presented in this paper aims to provide insights for future research addressing these challenges. Additionally, while this study examines four fundamental types of visual variations, it does not exhaust the infinite possibilities of visual transformations, leaving some long-tail cases unexplored.
@@ Appendix A Discussion
The benchmark’s component-level analysis offers fine-grained diagnostic information, guiding researchers toward targeted improvements. A key future direction lies in enhancing the multimodal alignment mechanisms. Additionally, the identified vulnerabilities are not merely the result of limited data or training strategies but also stem from fundamental architectural constraints requiring further investigation.
@@ Appendix B Prompt Template
Figure 7: Linear probing results for Vision Encoder showing position bias patterns similar to those observed in LVLMs.
@@ C.3 Object and Orientation Dataset
The object and orientation dataset is designed to evaluate model understanding of object positioning and directionality. We carefully select ten distinct object categories, including eight animals (shiba dog, cat, bear, eagle, snake, panda, turtle, and fish) and two vehicles (car and plane). This diverse selection allows us to assess model performance across varying object morphologies and complexities.

### v3

# Unveiling the Lack of LVLM Robustness to Fundamental Visual Variations: Why and Path Forward

@@ Abstract
Large Vision Language Models (LVLMs) excel in various vision-language tasks. Yet, their robustness to visual variations in position, scale, orientation, and context that objects in natural scenes inevitably exhibit due to changes in viewpoint and environment remains largely underexplored. To bridge this gap, we introduce V2R-Bench, a comprehensive benchmark framework for evaluating Visual Variation Robustness of LVLMs, which encompasses automated evaluation dataset generation and principled metrics for thorough robustness assessment. Through extensive evaluation on 21 LVLMs, we reveal a surprising vulnerability to visual variations, in which even advanced models that excel at complex vision-language tasks significantly underperform on simple tasks such as object recognition. Interestingly, these models exhibit a distinct visual position bias that contradicts theories of effective receptive fields, and demonstrate a human-like visual acuity threshold. To identify the source of these vulnerabilities, we present a systematic framework for component-level analysis, featuring a novel visualization approach for aligned visual features. Results show that these vulnerabilities stem from error accumulation in the pipeline architecture and inadequate multimodal alignment. Complementary experiments with synthetic data further demonstrate that these limitations are fundamentally architectural deficiencies, scoring the need for architectural innovations in future LVLM designs. [^1] Our code and data is available at https://github.com/toward-agi/Visual-Variations-Robustness.
[…]
Unveiling the Lack of LVLM Robustness to Fundamental Visual Variations:
Why and Path Forward
Zhiyuan Fanthanks: Equal contribution.     Yumeng Wang     Sandeep Polisetty     Yi R. (May) Fung Hong Kong University of Science and Technology zhiyuan.fan@connect.ust.hk    yrfung@cse.ust.hk
Figure 1: Our V2R-Bench benchmark contains: 1) An automated data generation pipeline that adds visual variations to existing benchmarks; (2) Synthetic data generation for the fundamental visual tasks proposed (e.g., object, direction, coordinate, path and OCR tasks). The subfigure on the right shows metrics for robustness testing, including performance consistency, semantic and token-level stability, and LLM-as-a-judge.
@@ 2 Introduction
While current benchmarks extensively evaluate models using images collected in specific scenarios, they overlook a more fundamental generalization capability: the robustness of LVLMs to visual variations. In input images, objects naturally exhibit diverse variations: spatial positions shift with changes in camera angles and viewpoints; object scales vary depending on viewing distances; orientations deviate from standard poses through rotations and inversions; and objects appear in a range of visual semantic contexts. These visual variations raise concerns about LVLMs’ robustness: whether models maintain consistent perception capabilities across all spatial positions in input images; what visual acuity threshold determines reliable performance; and how changes in orientation and visual context influence model behavior. Despite the importance of such robustness, current research has primarily focused on model robustness to socio-cultural factors (Ananthram et al., 2024), adversarial prompt attacks (Wu et al., 2024b; Liu et al., 2024b), or corrupted image inputs (Liu et al., 2024a), leaving the impact of natural visual variations largely unexplored.
@@ 3 Related Work
The rapid progress of LVLMs is driven by continuous advances in model architectures and the exponential growth of training datasets. On the architectural side, improvements in foundation language models (OpenAI et al., 2024; Touvron et al., 2023; Jiang et al., 2023; Bai et al., 2023a) and vision models (Radford et al., 2021; Oquab et al., 2024), combined with increasingly sophisticated multimodal alignment modules (Liu et al., 2023; Li et al., 2023b; Alayrac et al., 2022; Zhu et al., 2023; Wang et al., 2024b; Zhou et al., 2024), have established the core capabilities of LVLMs in cross-modal understanding and reasoning. Building on these foundations, the introduction of high-quality datasets (Chen et al., 2023a; Zhao et al., 2023; Wang et al., 2023b; Li et al., 2023c) for different training stages (e.g., multimodal alignment, instruction tuning, preference alignment) has enabled LVLMs to achieve strong performance across diverse real-world scenarios. Despite effectively leveraging pretrained vision and language components, the pipeline architecture can accumulate errors across modules. No systematic investigation has yet been conducted to attribute model failures to specific components, limiting interpretability and understanding of modality alignment.
@@
4 V2R-Bench Evaluation Framework
In this section, we propose a novel framework for automatically automatically generating diverse visual variations across task settings (4.1), along with a detailed description of the construction and characteristics of our evaluation dataset (4.2) and our evaluation protocols (4.3), as illustrated in Figure 1.
@@ 4.1 Automated Variations Generation
The automated data generation pipeline incorporates four fundamental visual variations that are prevalent in real-world scenarios. Position variations investigate whether LVLMs exhibit blind spots in their visual processing, where models may fail to accurately perceive visual information at certain spatial locations. Meanwhile, scale variations examine the perceptual boundaries of LVLMs when processing objects at different scales, similar to clinical vision tests for humans. Orientation variations challenge the ability of LVLMs to process objects at different rotational angles, which is crucial for real-world scenarios like robotic navigation where objects rarely appear in canonical orientations. Context variations test whether model predictions remain consistent across diverse environmental settings, to reveal whether LVLMs perform genuine visual perception or rely primarily on contextual cues for inference.
@@
where $P=\{1,...,W\}\times\{1,...,H\}$, $S=[s_{min},s_{max}]$, $R=[0,2\pi]$, and $C\in\mathcal{B}$ represent the sets of position, scale, rotation, and context variations respectively. For each original question-image pair, this generation process produces $|P|\times|S|\times|R|\times|C|$ variants in total, enabling a holistic exploration of the entire variation space.
4.2 Dataset Construction
The proposed automated generation pipeline is implemented across two categories of tasks.
The first category extends to existing multimodal benchmarks, focusing on scenarios where variations in position, scale, orientation, and context preserve ground-truth validity, ensuring that any performance changes reflect model robustness rather than ground-truth alteration.
The second category examines fundamental visual capabilities through object recognition and direction detection tasks (i.e. basic visual tasks), with target objects and directional indicators being systematically transformed through our visual variations using image processing algorithms and inpainting diffusion models (Corneanu et al., 2024; Lugmayr et al., 2022). The object dataset assesses a model’s understanding of object positioning and directionality, which is a 90-class object classification task. For the orientation dataset, the model performs an 8-class classification to detect an arrow’s direction for evaluating its basic spatial reasoning.
The final evaluation datasets contain a total of 428K images. Each category serves a distinct yet complementary purpose in our evaluation: the basic tasks provide controlled, interpretable measures of fundamental capabilities, while the extended benchmarks assess robustness in more naturalistic settings. The detailed implementation of these generation algorithms is provided in Appendix C.
@@ 4.3 Evaluation Metrics
Furthermore, LLM-as-a-judge (Zheng et al., 2023) is employed to emulate human assessment of LVLM-generated outputs under structured visual variations, providing an additional qualitative perspective on model robustness. The overall robustness is then determined by a weighted aggregation of these three evaluation dimensions.
@@
Table 1: Comprehensive evaluation of model performance (i.e., accuracy and aggregated robustness) under different visual variations, for object and direction detection tasks) with metrics defined in Section 4.3.
@@ 5.1 Component-level Analysis
$v=E_{v}(I;\theta_{v})\in\mathbb{R}^{N_{v}\times D_{v}}$ (5)
@@
$h=P(v;\theta_{p})\in\mathbb{R}^{N_{v}\times D_{h}}$ (6)
@@
The multimodal projector $P$ acts as a bridge between the visual feature and language embedding spaces, which raises two questions for its performance analysis: (RQ1) Do the projected features $h$ preserve the visual information contained in $v$, and (RQ2) Do the projected features $h$ align well with the embedding space of the language model $M$? To answer these questions, two analytical approaches are employed: (1) comparing the performance of pre-projection features $v$ and post-projection features $h$ on visual tasks to quantify potential visual information loss during multimodal alignment, and (2) measuring the discriminability between projected visual features $h$ and the language embeddings of their corresponding captions to assess the quality of modality alignment.
@@ 5.2 Visual-Linguistic Feature Analysis
$t=\text{topk}\left(\text{softmax}\left(hE^{\top}\right)\right),$ (8)
@@
Through these decoded language tokens, the semantic meaning captured by the aligned visual features becomes intuitively understandable, which helps address the following questions: (RQ3) how do language models interpret these aligned visual features? (RQ4) How robust is the semantics of aligned visual features to visual variations? These questions are investigated in detail in Section  6.2.
@@ 6.1 Impact of Visual Variations
The evaluation results across different LVLM architectures and model scales are presented in Table 1. Despite their impressive performance demonstrated on complex visual tasks, these models exhibit surprising vulnerability to simple visual variations, resulting in significantly degraded performance across basic visual tasks. Even proprietary models such as GPT-4o, Claude, and Gemini not only produce incorrect outputs, but also occasionally claim an inability to perceive the visual content entirely, as illustrated in Appendix F. Interestingly, despite claims that some carefully distilled smaller models outperform their larger counterparts on mainstream benchmarks Liu et al. (2024c); Bai et al. (2023b), our analysis reveals that scaling laws still hold for robustness: within the same model architecture, larger models consistently demonstrate better stability across visual variations.
@@
Table 2: Quantitative analysis of position bias on the proposed basic visual tasks: object detection and direction recognition. Experimental results show that these models consistently demonstrate better perception of surrounding regions compared to the central areas.
@@ 6.2 Component Analysis
Table 4: Linear probing accuracy (%) of vision language models on Object and Direction tasks reveals the impact of vision encoders/multimodal projectors in LVLMs. Notably, Qwen2-VL, which employs a native vision encoder trained directly with the language model rather than pretraining on image modality first and then aligning to language modality, shows no performance loss and even improvement after modality projection, suggesting the potential of unified architectures.
@@
Table 5: Accuracy of LVLM’s language model backbones on text-based tasks: comparing scenarios with tokenized background represented as asterisks (w/ BG) versus random words (w/o BG).
@@ 8 Conclusion
We present an evaluation framework, V2R-Bench, designed to assess the robustness of LVLMs against visual variations. Our results show significant vulnerabilities in existing LVLMs and identify their origins as twofold: insufficient multimodal alignment and error accumulation inherent in pipeline model architectures. While synthetic data augmentation Qin et al. (2025) seems as a mitigation strategy, fundamental advancement in the field requires a shift towards native multimodal architectures rather than the current approach of concatenating separate language and vision modalities. We aim to draw greater attention to the importance of LVLM visual robustness and inspire future research toward more robust architectural paradigms.
@@
Aligned with the paper track’s focus, this work primarily concentrates on identifying a novel problem, establishing corresponding evaluation methods, and providing initial analytical insights into these vulnerabilities. Regarding solutions to the identified vulnerabilities, only two categories of data construction approaches were explored in the automated construction pipeline, while potential improvements through architectural modifications, pre-training strategies, test-time self-correction or self-consistency approaches He et al. (2024); Wang et al. (2025), and novel modality alignment methods remain unexplored due to their substantial resource requirements. The detailed analysis presented in this paper aims to provide insights for future research addressing these challenges. Additionally, while this study examines four fundamental types of visual variations, it does not exhaust the infinite possibilities of visual transformations, leaving some long-tail cases unexplored.
@@ Appendix A Discussion
The benchmark’s component-level analysis offers fine-grained diagnostic information, guiding researchers toward targeted improvements. A key future direction lies in enhancing the multimodal alignment mechanisms. Additionally, the identified vulnerabilities are not merely the result of limited data or training strategies but also stem from fundamental architectural constraints requiring further investigation. Moreover, in the longer horizon, we believe that new techniques for enhancing LVLM visual variation robustness can help address the broader objectives of multimodal knowledge boundary awareness and hallucination mitigation as well He et al. (2025).
@@ Appendix B Prompt Template
Figure 7: Linear probing results for Vision Encoder show position bias patterns similar to those observed in LVLMs.
@@ C.3 Object and Orientation Dataset
The object dataset evaluates a model’s understanding of object positioning and directionality. It comprises ten carefully selected object categories: eight animals (shiba dog, cat, bear, eagle, snake, panda, turtle, and fish) and two vehicles (car and plane). This diverse set enables assessment of model performance across varied object morphologies and complexities. In the linear probing experiment, the model performs a 90-class classification task based on these objects.
In the orientation dataset, the model is instructed to detect the directions of the arrow in an image. There are in total 8 possible directions, including up, upper-right, right, bottom-right, bottom, bottom-left, left and upper-left. This comprehensive set of orientations ensures a robust assessment of the model’s capacity for basic spatial reasoning and its interpretation of symbolic directional cues. In the linear probing setting, the model performs a 8-class classification task based on these objects.